# XY-Cut++: Advanced Layout Ordering via Hierarchical Mask Mechanism on a Novel Benchmark

### v2

@@ 5 Experiments
BLEU-4 $\uparrow$ ARD $\downarrow$ Tau $\uparrow$ BLEU-4 $\uparrow$ ARD $\downarrow$ Tau $\uparrow$ BLEU-4 $\uparrow$ ARD $\downarrow$ Tau $\uparrow$
@@
Method Mask Mask Cross-Layout Pre-Cut Adaptive Scheme $\phi_{1}$ $\phi_{2}$ $\phi_{3}$ $\phi_{4}$ Dynamic Weights Multi-Stage BLEU-4 $\uparrow$ Baseline 0.797 +Pre-Mask 0.822 +MGS 0.905 0.914 0.923 0.962 +CMM 0.963 0.765 0.858 0.985 0.881 0.694 0.988
@@
Table 5: Reading Order Recovery Performance on Textual Content of OmniDocBench (Excluding Figures/Tables with Insignificant Layout Impact). Metric Key: BLEU-4 $\uparrow$ / ARD $\downarrow$ / Tau $\uparrow$. The red color indicates the best results and the blue color indicates the second-best results.
@@
Table 6: Model Efficiency and Semantic Information Usage on DocBench-100 and OmniDocBench. Key Metrics: FPS (Total Pages/Total Times) $\uparrow$. FPS values are averaged over 10 runs on Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz with 256GB memory. The red color indicates the best results and the blue color indicates the second-best results.
@@ 5.3.1 Progressive Component Ablation
  • XY-Cut Baseline: Achieves 0.749 BLEU-4 on complex layouts ($D_{c}$), showing limitations in handling complex-layout(e. g. L-shape) elements.
  • +Pre-Mask: Improves BLEU-4 by 6.9 points (from 0.749 to 0.818) on $D_{c}$ via adaptive thresholding (Equation (1), $\beta=1.3$), reducing false splits by 15.9%.
  • +MGS: Delivers a 19.7 absolute BLEU-4 gain on $D_{c}$ through three-phase segmentation, with density-aware splitting (Equation (5)) reducing ARD by 29.7%.
  • +CMM: Achieves near-perfect alignment (0.995 $\tau$) on $D_{c}$ through geometric constraints (Equations (7)-(10)), finalizing a 90.1% ARD reduction from baseline.
The complete model reduces ARD by 93.5% compared to baseline (0.139→0.009), demonstrating superior ranking consistency. Notably, our method maintains balanced performance across both subsets (0.988 $\mu$-BLEU), proving effective for diverse layout types.
@@ 5.3.2 Architectural Analysis
Pre-Mask Processing (Pre-Mask): To alleviate the ”L-shaped” problem, we applied preliminary masking on highly dynamic elements such as titles, tables, and figures (see Methodology Section 4.1). This approach reduced visual noise and improved reading order recovery, resulting in a 2.5-point BLEU-4 score increase, as shown in Table 3.
extbfMulti-Granularity Segmentation (MGS): Mask cross-layout is a very direct method to solve the problem that XY-Cut cannot segment L-shaped input. Pre-Cut realizes preliminary sub-page division through page analysis, thereby avoiding page content mixing affecting sorting. The adaptive splitting strategy enables reasonable segmentation through real-time density estimation. Table 3 shows additive benefits of the mask cross-layout (+8.3), Pre-Cut (+9.2), and the adaptive splitting scheme (+10.1) over the baseline (0.822 BLEU-4).
@@ 5.3.3 Benchmark Comparison
As shown in Table 4, our approach establishes new benchmarks across all evaluation dimensions. Specifically, it outperforms XY-Cut by a significant margin, achieving a +23.7 absolute improvement on $D_{c}$ (from 74.9 to 98.6). Additionally, it surpasses LayoutReader by +5.3 on $D_{r}$ (from 94.6 to 98.9), despite not using any learning-based components. Furthermore, our method achieves a Kendall’s $\tau$ of 0.996 overall, indicating near-perfect ordinal consistency (with $p<0.001$ in the Wilcoxon signed-rank test). Visual results presented in Fig. 9 and 9 further demonstrate the robustness in handling multi-column layouts and cross-page elements, where previous methods frequently fail.
@@ 6 Conclusion
In this work, we address the critical challenges in document layout recovery and semantic conversion, particularly focusing on the needs of Retrieval-Augmented Generation (RAG) and Large Language Models (LLMs) for accurate information extraction. We introduce a hierarchical document parsing framework that incorporates a Mask mechanism-based layout recovery algorithm. This framework enables precise reading order recovery through pre-mask processing, multi-granularity segmentation, and cross-modal matching. Additionally, it achieves superior performance on the DocBench-100 and OmniDocBench datasets, outperforming baseline methods and providing robust support for downstream RAG and LLM tasks. It is noteworthy that while traditional methods are known for their computational efficiency, they often fall short in performance. Our approach maintains simplicity and efficiency while matching or surpassing deep learning methods. This advantage stems from two innovations: (1) shallow semantic labels used as structural priors, and (2) a hierarchical mask mechanism that captures document topology. These improve performance without sacrificing efficiency and suggest promising directions for future work. Code will be released at: https://github.com/liushuai35/PaddleXrc.git.
7 Limitations and Future Work
Our method currently struggles to handle pages with nested semantic structures, like sub-pages within a single page, as it cannot effectively sort elements within these complex layouts. To tackle this, we plan to enhance our framework by adding a sub-page detection module. This module will identify sub-page regions, enabling element sorting within them while maintaining the natural reading order (top-to-bottom, left-to-right) for the sub-pages themselves. By balancing global and local sorting, this approach aims to enhance the robustness of our method for complex page structures.
8 Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v3

@@ 5 Experiments
BLEU-4 ARD Tau BLEU-4 ARD Tau BLEU-4 ARD Tau
@@
Method Mask Mask Cross-Layout Pre-Cut Adaptive Scheme $\phi_{1}$ $\phi_{2}$ $\phi_{3}$ $\phi_{4}$ Dynamic Weights Multi-Stage BLEU-4 Baseline 0.797 +Pre-Mask 0.822 +MGS 0.905 0.914 0.923 0.962 +CMM 0.963 0.765 0.858 0.985 0.881 0.694 0.988
@@
Table 5: Reading Order Recovery Performance on Textual Content of OmniDocBench (Excluding Figures/Tables with Insignificant Layout Impact). Metric Key: BLEU-4 $\uparrow$ / ARD $\downarrow$ / Tau $\uparrow$. Best results are in bold; second-best are underlined.
@@
Table 6: Model Efficiency and Semantic Information Usage on DocBench-100 and OmniDocBench. Key Metrics: FPS (Total Pages/Total Times) $\uparrow$. FPS values are averaged over 10 runs on Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz with 256GB memory. Best results are in bold; second-best are underlined.
@@ 5.3.1 Progressive Component Ablation
  • XY-Cut Baseline: Achieves 0.749 BLEU-4 on complex layouts ($D_{c}$), showing limitations in handling complex layouts (e.g., L-shaped) elements.
  • +Pre-Mask: Improves BLEU-4 by 6.9 points (from 0.749 to 0.818) on $D_{c}$ via adaptive thresholding (Eq. (1), $\beta=1.3$), reducing false splits by 15.9%.
  • +MGS: Delivers a 19.7 absolute BLEU-4 gain on $D_{c}$ through three-phase segmentation, with density-aware splitting (Eq. (5)) reducing ARD by 29.7%.
  • +CMM: Achieves near-perfect alignment (0.995 $\tau$) on $D_{c}$ through geometric constraints (Eq. (9)–(12)), finalizing a 90.1% ARD reduction from baseline.
The complete model reduces ARD by 93.5% compared to baseline (0.139 $\rightarrow$ 0.009), demonstrating superior ranking consistency. Notably, our method maintains balanced performance across both subsets (0.988 $\mu$-BLEU), proving effective for diverse layout types.
@@ 5.3.2 Architectural Analysis
Pre-Mask Processing (Pre-Mask): To alleviate the ”L-shaped” problem, we applied preliminary masking on highly dynamic elements such as titles, tables, and figures (see Methods Section 4.1). This approach reduced visual noise and improved reading order recovery, resulting in a 2.5-point BLEU-4 score increase, as shown in Table 3.
Multi-Granularity Segmentation (MGS): Mask cross-layout is a very direct method to solve the problem that XY-Cut cannot segment L-shaped input. Pre-Cut realizes preliminary sub-page division through page analysis, thereby avoiding page content mixing affecting sorting. The adaptive splitting strategy enables reasonable segmentation through real-time density estimation. Table 3 shows additive benefits of the mask cross-layout (+8.3), Pre-Cut (+9.2), and the adaptive splitting scheme (+10.1) over the baseline (0.822 BLEU-4).
@@ 5.3.3 Benchmark Comparison
As shown in Table 4, our approach establishes new benchmarks across all evaluation dimensions. Specifically, it outperforms XY-Cut by a significant margin, achieving a +23.7 absolute improvement on $D_{c}$ (from 74.9 to 98.6). Additionally, it surpasses LayoutReader by +5.3 on $D_{r}$ (from 94.6 to 98.9), despite not using any learning-based components. Furthermore, our method achieves a Kendall’s $\tau$ of 0.996 overall, indicating near-perfect ordinal consistency (with $p<0.001$ in the Wilcoxon signed-rank test). Visual results presented in Figure 9 and 9 further demonstrate the robustness in handling multi-column layouts and cross-page elements, where previous methods frequently fail.
@@
6 Discussion
XY-Cut++ closes the gap between classical projection-based methods and neural models for block-level reading-order recovery through a simple three-stage pipeline. Directional ordering (left-to-right, top-to-bottom) is largely solved; the remaining pain points are fine-grained semantic segmentation (e.g., ambiguous block boundaries and caption linkage; see Figure 17) and sub-page structures. Across DocBench-100 and OmniDocBench, XY-Cut++ delivers large, consistent gains (Table 2, Table 5) while retaining high throughput on CPU (Table 6). On $D_{c}$, BLEU-4 rises from 0.749 to 0.986 and ARD falls from 0.233 to 0.023, with visual results (Figures 9,9) confirming robustness on multi-column and L-shaped cases.
Mechanistically, the three components are complementary: Pre-Mask suppresses dominant elements to expose a stable backbone; MGS uses an adaptive density $\tau_{d}$ (Eq. (4)) to choose the split axis (Eq. (5)); and CMM applies four geometric constraints (Eqs. (9)–(12)) with scale-aware (Eq. (13)) and semantic-specific weights (Eq. 14). Ablations indicate edge-weighted margins are especially effective.
Practically, reliable block-level ordering benefits RAG/LLM preprocessing by improving chunking and caption attachment, and XY-Cut++ is easy to deploy after a detector with coarse labels (e.g., PP-DocLayout). DocBench-100 helps standardize block-level evaluation and emphasizes difficult page topologies.
Limitations primarily stem from label granularity for fine-grained segmentation and from nested “sub-pages” (Figure 17); detector quality also matters. Heuristic hyperparameters (e.g., $\beta$ in Eq. (1), $\theta_{v}$ in Eq. (5)) may require domain tuning. Future work targets: (i) sub-page detection with hierarchical reasoning to localize ordering, (ii) learning the split policy and edge weights from weak supervision to improve within-block segmentation, and (iii) coupling lightweight language priors (caption/title cues) with end-to-end RAG/LLM evaluation.
@@ 7 Conclusion
We studied block-level reading-order recovery for real-world documents, a prerequisite for reliable RAG/LLM pipelines. We proposed XY-Cut++, a hierarchical, geometry-aware framework that integrates pre-mask processing, multi-granularity segmentation, and cross-modal matching. XY-Cut++ delivers precise block ordering and achieves state-of-the-art performance on DocBench-100 and OmniDocBench, while maintaining high throughput on CPU (cf. Tables 26).
This accuracy–efficiency balance is achieved via two design choices: (i) shallow semantic labels used as structural priors and (ii) a hierarchical mask mechanism that stabilizes XY-Cut through density-aware splitting and edge-weighted matching. These components jointly improve accuracy without sacrificing throughput, making XY-Cut++ a practical module for production-grade document parsing. Beyond the algorithm, DocBench-100 offers a focused, block-level benchmark that emphasizes challenging page topologies and promotes standardized evaluation. Future work will extend XY-Cut++ with sub-page detection and lightweight language priors to better handle nested and fine-grained structures (see Section 6).